OpenStudy (anonymous):

find f'(a) f(x)=x^-2

OpenStudy (anonymous):

\[f(x)=x^n\] \[f'(x)=nx^{n-1}\]

OpenStudy (anonymous):

or you can do it the long way \[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

i need to somehow get it into f(a+h)-f(a)/h

OpenStudy (anonymous):

ya were learning the long way

OpenStudy (anonymous):

\[f(x)=x^{-2}\] \[f(y)=y^{-2}\] \[f(x+100)=(x+100)^{-2}\] \[f(1)=1^{-2}\] \[f(x+h)=???\]

OpenStudy (anonymous):

(a+h)^-2 - (a)^-2 --------------- h

OpenStudy (anonymous):

im blanking on negative exponents, let me check my notes

OpenStudy (anonymous):

right now lets just ignore the h for now we'll get back to that \[(a+h)^{-2} - (a)^{-2}=\frac{1}{(a+h)^{2}} - \frac{1}{(a)^2}\]

OpenStudy (anonymous):

now just simplify that equation, then put it all over h

OpenStudy (anonymous):

ohhhhh your the man/woman

OpenStudy (anonymous):

i prefer idiot

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks a lot, that really made things clear

OpenStudy (anonymous):

hey idiot, i got a weird answer after working through it

OpenStudy (anonymous):



OpenStudy (anonymous):

f(x) = x^-2 let f(x) = y y=x^-2 y=1/x^2 x^2=1/y x=\[\sqrt{(1/y)}\] therefore f'(x)=\[\sqrt{(1/x)}\]

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] \[f(x)=x^{-2}\] \[\lim_{h \rightarrow 0}\frac{(x+h)^{-2}-x^{-2}}{h}\] \[\lim_{h \rightarrow 0}\frac{\frac{1}{(x+h)^{2}}-\frac{1}{x^{2}}}{h}\] \[\lim_{h \rightarrow 0}\frac{\frac{1}{x^2+2xh+h^2}-\frac{1}{x^{2}}}{h}\] \[\lim_{h \rightarrow 0}\frac{\frac{x^2}{x^2(x^2+2xh+h^2)}-\frac{x^2+2xh+h^2}{x^{2}(x^2+2xh+h^2)}}{h}\] \[\lim_{h \rightarrow 0}\frac{\frac{-2xh-h^2}{x^{2}(x^2+2xh+h^2)}}{h}\] \[\lim_{h \rightarrow 0}\frac{-2x-h}{x^{2}(x^2+2xh+h^2)}\] \[\frac{-2x}{x^2*x^2}\] \[\frac{-2x}{x^4}\] \[\frac{-2}{x^3}\] \[f'(x)=-2x^{-3}\]

OpenStudy (anonymous):

you can only cross multiply only if the 2 fractions are equal to each other in this case you can only find the GCD or LCD whichever it is and then combine it, kind of like cross multiplying